Yellow Corp. is evaluating an extra dividend versus a share repurchase. In either case, $5,500 would be spent. Current earnings are $1.16 per share and the stock currently sells for $45 per share. There are 2,300 shares outstanding. Ignore taxes and other imperfections.

NOTE: Fractional shares are possible (Ex. 0.54 shares)

Dividend =   2.39

Shares outstanding = 2300

Stock price = 42.61

After the $2.39 dividend, the price falls to $42.61 per share. What are earnings per share (EPS) and the price earnings (P/E) ratio? Enter your answers rounded to 2 DECIMAL PLACES.

EPS=

P/E Ratio =
Answer #1

Solution:

If the company choses to pay extra dividend then it means shares are not repurchased. Hence, there will be no impact on EPS, it will remain same $1.16 per share.

  • Revised Market Price per share (given) = $42.61
  • P/E Ratio = \frac{Market Price of the share}{EPS}

=42.61/1.16

= 36.73